write an algebraic expression for each of the following
a)suman is 7 years younger to his brother aman
b)diameter of a circle is twice the radius
c)successor of a natural number n

Found 2 solutions by Theo, ikleyn:
Answer by Theo(13342)

a)suman is 7 years younger to his brother aman

x = y - 7

x = suman's age
y = aman's age

b)diameter of a circle is twice the radius

d = 2 * r

d = diameter
r = radius

c)successor of a natural number n

s = n + 1

s is the successor to the natural number n.
n is the natural number.








Answer by ikleyn(52781)

.
write an algebraic expression for each of the following

a) suman is 7 years younger to his brother aman    S = A - 7 years old.

b) diameter of a circle is twice the radius        D = 2R.

c) successor of a natural number n                 n = n+1, where n is natural number.